\begin{thebibliography}{10}

\bibitem{BKMTV}
E.~A. Bespalov, D.~S. Krotov, A.~A. Matiushev, A.~A. Taranenko, and K.~V.
  Vorob'ev.
\newblock Perfect $2$-colorings of {H}amming graphs.
\newblock {\em
  \href{http://onlinelibrary.wiley.com/journal/10.1002/(ISSN)1520-6610}{J.
  Comb. Des.}}, 29(6):367--396, June 2021.
\newblock \DOI{10.1002/jcd.21771}.

\bibitem{QextNewEdition}
I.~Bouyukliev and S.~Bouyuklieva.
\newblock {\texttt{QextNewEdition}}.
\newblock
  \url{http://www.moi.math.bas.bg/moiuser/~data/Software/QextNewEdition}.

\bibitem{BFWW2006}
I.~Bouyukliev, V.~Fack, W.~Willems, and J.~Winne.
\newblock Projective two-weight codes with small parameters and their
  corresponding graphs.
\newblock {\em \href{http://link.springer.com/journal/10623}{Des. Codes
  Cryptography}}, 41(1):59--78, 2006.
\newblock \DOI{10.1007/s10623-006-0019-1}.

\bibitem{BvM:SRG}
A.~E. Brouwer and H.~Van~Maldeghem.
\newblock {\em Strongly Regular Graphs}, volume 182 of {\em Encycl. Math.
  Appl.}
\newblock Cambridge: Cambridge University Press, 2022.

\bibitem{FDF:PerfCol}
D.~G. Fon-Der-Flaass.
\newblock Perfect $2$-colorings of a hypercube.
\newblock {\em \href{http://link.springer.com/journal/11202}{Sib. Math. J.}},
  48(4):740--745, 2007.
\newblock \DOI{10.1007/s11202-007-0075-4} translated from
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=smj\&option_lang=eng}{Sib.
  Mat. Zh.} 48(4) (2007), 923-930.

\bibitem{GAMS}
{\em GAMS Development Corporation. General Algebraic Modeling System (GAMS)
  Release 24.2.1.}
\newblock Washington, DC, USA, 2013.

\bibitem{HamHel96}
N.~Hamada and T.~Helleseth.
\newblock A characterization of some {$\{3v_2+v_3,3v_1+v_2;3,3\}$}mminihypers
  and some {$[15,4,9;3]$}-codes with {$B_ 2=0$}.
\newblock {\em \href{http://www.sciencedirect.com/science/journal/03783758}{J.
  Stat. Plann. Inference}}, 56(1):129--146, 1996.
\newblock \DOI{10.1016/S0378-3758(96)00014-6}.

\bibitem{KooKroMar:Ch7tables}
J.~H. Koolen, V.~N. Potapov, and W.~J. Martin.
\newblock Completely regular codes: Tables of small parameters for binary and
  ternary {H}amming graphs.
\newblock In Shi and Sol\'e \cite{CRCinDRG}, chapter~7.
\url{https://www.researchgate.net/publication/385378754_Completely_Regular_Codes_Tables_of_Small_Parameters_for_Binary_and_Ternary_Hamming_Graphs}

\bibitem{KroPot:Ch1}
D.~S. Krotov and V.~N. Potapov.
\newblock Completely regular codes and equitable partitions.
\newblock In Shi and Sol\'e \cite{CRCinDRG}, chapter~1.
\url{https://www.researchgate.net/publication/382042905_Completely_Regular_Codes_and_Equitable_Partitions}

\bibitem{CRCinDRG}
M.~Shi and P.~Sol\'e, editors.
\newblock {\em Completely Regular Codes in Distance Regular Graphs}.
\newblock Chapman \& Hall/CRC Monographs and Research Notes in Mathematics. CRC
  Press, Boca Raton FL, 2025.

\end{thebibliography}

\documentclass{article}
\usepackage{amsfonts,amsmath,amsthm,amssymb}
\usepackage{xcolor,graphicx}
\usepackage{comment}
\usepackage{hyperref}

\title{Completely regular codes in graphs covered by a Hamming graph\thanks{The work was supported by the grant of The Natural Science Foundation of Hebei Province (project No. A2023205045)}}
\author{Sergey Goryainov\thanks{School of Mathematical Sciences, Hebei International Joint Research Center for Mathematics and Interdisciplinary Science,
Hebei Key Laboratory of Computational Mathematics and Applications, Hebei Workstation for Foreign Academicians,
Hebei Normal University, Shijiazhuang 050024, P.R. China}\ and Denis Krotov\thanks{School of Mathematical Sciences, Hebei Key Laboratory of Computational Mathematics
and Applications, Hebei Normal University, Shijiazhuang 050024, P. R. China}
\thanks{Sobolev Institute of Mathematics, 
Novosibirsk 630090, Russia}}
\date{}

\newcommand{\FF}{\mathbb{F}}
\newcommand{\ZZ}{\mathbb{Z}}
\newcommand{\Aut}{\mathrm{Aut}}
\newcommand{\Cay}{\mathrm{Cay}}
\newcommand{\vx}{}

\newtheorem{proposition}{Proposition}
\newtheorem{lemma}{Lemma}

\newtheorem{corollary}{Corollary}
\theoremstyle{remark}
\newtheorem{remark}{Remark}
\theoremstyle{definition}
\newtheorem{definition}{Definition}

\begin{document}
\maketitle
\begin{abstract}
In Cayley graphs on the additive group of a small vector space over GF$(q)$, $q=2,3$, we look for
completely regular (CR) codes whose parameters are new in Hamming graphs over the same field. The existence of a CR code in such Cayley graph $G$ implies the existence of a CR code with the same parameters in the corresponding Hamming graph that covers $G$. In such a way, we find several completely regular codes with new parameters in Hamming graphs over GF$(3)$. The most interesting findings are two new CR-$1$ (with covering radius~$1$) codes that are independent sets (such CR are equivalent to optimal orthogonal arrays attaining the Bierbrauer--Friedman bound) and one new CR-$2$. By recursive constructions, every knew CR code induces an infinite sequence of CR codes (in particular, optimal orthogonal arrays if the original code was CR-$1$ and independent).
In between, we classify feasible parameters of CR codes in several strongly regular graphs.
\end{abstract}

\section{Introduction}

The main result of the current report
is finding completely regular (CR) codes with new parameters
in Hamming graphs $H(n,3)$.
The summary with new parameters can be found in 
Section~\ref{s:sum}. 
They update the tables of small parameters of CR codes
in~\cite{KooKroMar:Ch7tables}. 
Most of new parameters update the smallest~$n$ for which
CR codes with given intersection array are known to exist in $H(n,3)$. However, the discovered intersection array $\{21,4;2,21\}$ is completely new;
 $\{21,4;2,21\}$-CR codes were not know before
 in $H(n,3)$ for any~$n$.

Covering-radius-$1$ CR codes that are independent sets in Hamming graphs corresponds to optimal orthogonal arrays attaining the Bierbrauer--Friedman bound, see e.g. \cite[Sect.\,1.4.3.1]{KroPot:Ch1}. 
We have found two new such codes. 
They are first such nonbinary codes whose cardinality
is not a power of the prime divisor of~$q$ (and so, codes with the same parameters cannot be constructed 
as linear or additive codes).
Because of recursive constructions described in Section~\ref{s:constr}, now we have an infinite series
of such codes.

\section{Preliminaries}

In our note, all graphs are simple, without loops or multiple edges. 
However, some introduced concepts and 
claims are valid for multigraphs,
see the details e.g. in \cite{KroPot:Ch1}.

As usual, $\FF(q)$ denotes the finite field
of prime power order $q$ 
and $\FF(q)^n$ denotes the vector space
of $n$-tuples ($n$-words) over~$\FF(q)$.
In this note, we mostly consider
the cases $q=2$ and $q=3$; 
in particular, Cayley graphs,
defined below,
are considered over the additive group of vector spaces over $\FF(q)$, $q=2,3$.

\newcommand{\Id}{\mathrm{Id}}
\begin{definition}[\bf\boldmath Cayley graph]
\label{1def:Cayley}
The \emph{Cayley graph} $\Cay(G,S)$ on a group $G$ (written in the multiplicative notation)
with \emph{connecting} set
$S$, $S\subset G\backslash\{\Id\}$, $S=S^{-1}$,
is the graph whose vertices
are the elements of~$G$ and whose edge multiset~$E$ is $\{\{x,ax\}:\ x\in G, a\in S\}$
(equivalently, $\{x,y\}\in E$ if and only if $y\cdot x^{-1} \in S$).
The Cayley graph is simple if and only if
$\Id \not\in S$ and $S$ is a simple set.
\end{definition}

\begin{definition}[\bf\boldmath 
Hamming graph
]
\label{1def:Hamming}
The \emph{complete graph} $K_q$ is the simple graph of order~$q$ where every two different vertices are adjacent.
The \emph{Hamming graph} $H(n,q)$ is the Cartesian product of $n$ copies of~$K_q$.
According to the definition of the Cartesian product, the vertex set of $H(n,q)$
consists of all $n$-words in some alphabet of order~$q$, which,
if $q$ is a prime power, is often associated with the elements of the Galois field
$\FF_q=\mathrm{GF}(q)$ of order~$q$.
In this case, the vertex set of $H(n,q)$ forms an $n$-dimensional vector space~$\FF_q^n$ over~$\FF_q$,
and $H(n,q)$ is a Cayley graph on its additive group~$\FF_q^{n+}$. The natural minimum-path distance in a Hamming graph is called the \emph{Hamming distance}.
\end{definition}

\begin{definition}[\bf\boldmath clique, coclique = independent set]
\label{1def:clique}
In a graph, a set of vertices that are mutually adjacent (mutually non-adjacent) to each other
is called a \emph{clique} (\emph{coclique}, or \emph{independent set}) or, more specifically,
\emph{$k$-clique} (\emph{$k$-coclique}) where $k$ is the cardinality of this set.
\end{definition}

\begin{definition}[\bf\boldmath perfect coloring, equitable partition, quotient matrix]
\label{1def:equitable}
A $k$-coloring of a graph~$\Gamma$ is called \emph{perfect}
if there is a $k$-by-$k$
matrix $(S_{i,j})_{i,j\in \Sigma}$
(the \emph{quotient} matrix)
whose rows and columns are indexes by colors,
such that every vertex of color~$i$
has exactly~$S_{i,j}$ neighbors
of color~$j$. An ordered partition $(C_0, \ldots, C_{k-1})$
of the vertex set of~$\Gamma$ into nonempty sets
$C_0$, \ldots, $C_{k-1}$ is called \emph{equitable},
or an \emph{equitable $k$-partition} if the vertex coloring~$f$
defined as $f(x)=i$ if $x\in C_i$ is perfect.
We will also use
the shorter forms
``$S$-perfect coloring'' for ``perfect coloring with quotient matrix~$S$''
and 
``equitable $S$-partition'' for ``equitable partition with quotient matrix~$S$''.
\end{definition}

\begin{definition}[\bf\boldmath covering radius, completely regular (CR) code, completely regular partition]\label{1d:crc}
\label{1def:crc}
Given a nonempty set~$C$ of vertices of a graph~$\Gamma=(V,E)$,
the \emph{distance partition} with respect to~$C$
is the partition $(C^{(0)},C^{(1)},\ldots ,C^{(\rho )})$ of~$V$
with respect to the distance from~$C$: $C^{(i)}=\{\vx x\in V:
\ d(\vx x,C)=i\}$. The number~$\rho=\rho(C)$ of
non-empty cells different from~$C$
in the distance partition
is called the
\emph{covering radius} of~$C$.
A nonempty set~$C$ (code) of vertices of a graph
is called
\emph{completely regular} (\emph{CR}) (completely regular set or completely regular code) 
if the partition of the vertex set
with respect to the distance
from~$C$ is equitable,
or, equivalently,
if the coloring $f(x)=d(x,C)$ of the vertex set with respect to the distance from~$C$
is perfect.
The quotient matrix
$(S_{i,j})_{i,j=0}^\rho$
corresponding
to a completely regular code
is always tridiagonal, and for regular graphs it is often written
in the form of the \emph{intersection array}
$$\{S_{0,1},S_{1,2},\ldots,S_{\rho-1,\rho}; S_{1,0},S_{2,1},\ldots,S_{\rho,\rho-1}\}$$
(the diagonal elements of the quotient matrix are uniquely determined from
the intersection array and the degree of the graph).

We will use the abbreviation ``$A$-CR'' for ``completely regular with intersection array~$A$''
or ``completely regular with quotient matrix~$A$'',
depending on the context, and 
the abbreviation ``CR-$\rho$'' for 
``completely regular with covering radius~$\rho$''.
\end{definition}

A very important special case
of perfect colorings
is graph coverings.

\begin{definition}[\bf\boldmath graph covering]\label{1d:gr-cov}
\label{1def:covering}
Assume that $f$ is a perfect coloring of a graph~$\Gamma$
with symmetric quotient matrix~$S$ and $\overline P$ is the equitable partition corresponding to~$f$.
Then $S$ can be considered as the adjacency matrix of
some graph, denoted $\Gamma/\overline P$, on the set of colors
(if $S$ is a $\{0,1\}$-matrix with zero diagonal,
then $T$ is a simple graph),
and we say that the coloring~$f$,
as well as the corresponding partition~$\overline P$,
is a \emph{covering} of the \emph{target} graph~$\Gamma/\overline P$
by a \emph{cover} graph~$\Gamma$.
In other words, a covering is a locally bijective
(bijective on the set of edges incident to a vertex)
surjective homomorphism from~$\Gamma$ to~$\Gamma/\overline P$.
\end{definition}

\begin{lemma}\label{1l:cay-cov}
    A surjective homomorphism $\varphi$ 
    from a group~$G$ onto a group $\varphi(G)$
    is a covering of $\Cay(\varphi(G),\varphi(S))$
    by $\Cay(G,S)$, for any multiset~$S$ of elements of~$G$ such that $S^{-1}=S$.
\end{lemma}

\begin{corollary}
    Any Cayley graph $\Cay(\FF_q^{k+},S)$ on the additive group of $\FF_q^{k}$, $q\in\{2,3\}$,
    is covered by $H(|S|/(q-1),q)$. Moreover, the corresponding covering is a homomorphism
    of $\FF_q^{n}$, $n=|S|/(q-1)$, onto $\FF_q^{k}$.
\end{corollary}

It is also known that 
\begin{lemma}\label{1l:conc}
The existence
    of an equitable $S$-partition of a graph $G$ implies
the existence of an equitable $S$-partition
of any graph~$G$ that covers~$G$.
\end{lemma}

By general algebraic arguments, it is easy to see that
Cayley graphs $\Cay(\FF_q^n,S)$, $q=2,3$, correspond
to vector subspaces (linear codes) of $\FF_q^n$, 
$(q-1)n=|S|$ with minimum Hamming distance between elements (codewords) at least~$3$.
By classifying small codes with the software
\cite{QextNewEdition}, we classify small Cayley graphs
with the corresponding parameters.

The corresponding between Cayley graphs and linear codes is as follows.
For a $k\times n$ matrix $H$ over $\FF_q$
without two collinear columns,
we denote by $G(H)$ the \emph{syndrome graph},
whose vertex set is $\FF_q^k$ and two vectors are adjacent if and only if their difference is collinear to a column of~$H$. Clearly, $G(H)$ is a Cayley graph
on $\FF_q^k$ where the connecting set is the set
of all nonzero vectors collinear to one of the columns.
On the other hand, the kernel of~$H$ is a 
vector subspace of $\FF_q^n$, a linear code $C$ with 
minimum Hamming distance at least~$3$ (distances~$1$ and $2$ are impossible because there are no two collinear columns in~$H$); the matrix~$H$ is known as a \emph{check} matrix of the code~$C$.

\section{Constructions of CR codes}
\label{s:constr}

\begin{proposition}\label{p:+t}
  If there is an equitable partition of $H(n,q)$ with 
quotient matrix $S$, then for every nonnegative integer~$t$ there is an equitable 
partition of $H(n+t,q)$ with 
quotient matrix $S+t(q-1)\mathrm{Id}$.
In particular, if there is a CR code with intersection array $A$ in $H(n,q)$, then $A$-CR code exists 
in $H(n+t,q)$.
\end{proposition}

\begin{proposition}\label{p:*s}
  If there is an equitable partition of $H(n,q)$ with 
quotient matrix~$S$, then for every positive integer~$s$ there is an equitable 
partition of $H(sn,q)$ with 
quotient matrix $sS$.
In particular, if there is a CR code with intersection array $A$ in $H(n,q)$, then $sA$-CR code exists 
in $H(sn,q)$.
\end{proposition}
\begin{proof}
If $(C_0,\ldots,C_k)$ is an equitable $S$-partition
of $H(n,q)$, then
$(C'_0,\ldots,C'_k)$,
where 
$$C'_i =  \{ (x_0,\ldots,x_{s-1})\in (\ZZ_q^n)^s:\ x_0+\ldots+x_{s-1} \in C_i \},$$
is an equitable $sS$-partition
of $H(sn,q)$. Indeed, if $x\in C_i$ has a neighbor
$x+e$ in $C_j$,
then every $(x_{0},\ldots,x_{s-1})\in C'_i$ such that
$y^{(l)}=x_0+\ldots+x_{s-1}=x$ has $s$ corresponding neighbors
$(x_1,\ldots,x_{l-1},x_l+e,x_{l+1},\ldots,x_s)$,
$l=0,...,s-1$ in~$C'_j$.
\end{proof}

\begin{proposition}[splitting {\cite[Th.\,4.13]{BKMTV}}]\label{p:split}
  If there is an equitable $[[a,b],[c,d]]$-partition $(C_0,C_1)$, $a\le c$, 
  in $H(n,q)$ ($a+c=b+d=(q-1)n$),
  where $q$ is a power of a prime,
  then in $H(qn+c-a,q)$
  there is an equitable $2$-partition 
  with quotient matrix 
$\begin{pmatrix}
    a+(i{-}1)c & (q{-}i)c + qb \\
    ic & qb+a+(q{-}i)c
\end{pmatrix}
$,
$i=1,...,q-1$.
\end{proposition}
\begin{proof}
We describe the construction for the case when
$q$ is prime. For a prime power case,
the construction is the same but one need to use
$\FF_q$ instead of $\ZZ_q$.
We start with constructing an equitable $\begin{pmatrix}qa&qb\\qc&qd\end{pmatrix}$-partition $(C'_0,C'_1)$ as in the proof of Proposition~\ref{p:*s}, where $s=q$.
Then, we split $C'_0$ into $q$ subcells
$C^{\prime 0}_0$, \ldots, $C^{\prime q-1}_0$,
according to the value of
$\alpha(y_0,\ldots,y_{q-1})=1\cdot|y_1|+ \ldots + (q-1)\cdot |y_{q-1}|$
for each $(y_0,\ldots,y_{q-1})$ in~$C'_0$.
It is not difficult to see that all $q$
vectors $y^{(l)}$, $l=0,...,s-1$ have distinct
values of $\alpha(y^{(l)})$.
It follows that any vertex of $H(qn,q)$ has equal numbers of neighbors in $C^{\prime 0}_0$, \ldots, $C^{\prime q-1}_0$, and
$(C^{\prime 0}_0, \ldots, C^{\prime q-1}_0,C^{\prime}_1)$ is an equitable partition with
quotient matrix 
$$
\begin{pmatrix}
    a & \cdots & a & qb \\
    \vdots & \ddots & \vdots & \vdots \\
    a & \cdots & a & qb \\
    c & \cdots & c & qd
\end{pmatrix}.
$$
Next, we construct a partition 
$(D^{ 0}_0, \ldots, D^{ q-1}_0,D_1)$ of $H(qn+1,q)$
as follows: 
$$ D^{ i}_0 =
\{(x,y) \in \ZZ_q^{qn+1}:\ 
y \in \ZZ_q,\ x\in C^{\prime i+y}_0\},$$
$$ D_0 =
\{(x,y) \in \ZZ_q^{qn+1}:\ 
y \in \ZZ_q,\ x\in C_0\}.$$
It is an equitable partition with quotient matrix
$$
\begin{pmatrix}
    a & a{+}1 & \cdots & a{+}1 & qb \\
    a{+}1 & a & \cdots & a{+}1 & qb \\
    \vdots & \vdots & \ddots & \vdots & \vdots \\
    a{+}1 & a{+}1 & \cdots & a & qb \\
    c & c &  \cdots & c & qd+q{-}1
\end{pmatrix}.
$$
Indeed, a vertex $x$ in $D^{ i}_0$ 
sees one neighbor in each $D^{j}_0$, $j\ne i$,
differing with $x$ in the last coordinate.

Repeating the last step $c-a$ times, we obtain
an equitable partition $(E^{0}_0, \ldots, E^{ q-1}_0,E_1)$
of $H(nq{+}c{-}a,q)$
with quotient matrix
$$
\begin{pmatrix}
    a & c & \cdots & c & qb \\
    c & a & \cdots & c & qb \\
    \vdots &\vdots &  \ddots & \vdots & \vdots \\
    c & c & \cdots & a & qb \\
    c & c &  \cdots & c & qd+(q{-}1)(c{-}a)
\end{pmatrix}.
$$
Now, because of equal coefficients 
in cells of the quotient matrix,
we can get an equitable $2$-partition
by unifying the first $i$ cells and
the remaining $q-i+1$ cells, $i=1,...,q-1$.
The resulting quotient matrix is
$$
\begin{pmatrix}
    a+(i{-}1)c & (q{-}i)c + qb \\
    ic & qb+a+(q{-}i)c
\end{pmatrix}.
$$
\end{proof}

\begin{remark}
The general splitting construction
(see \cite[Sect.\,3]{FDF:PerfCol} for $q=2$ and \cite[Th.\,4.14]{BKMTV} in general)
is more complicate and can sometimes get a better result
if the first cell $C_0$ of the original partition can be split into $s$-faces
(sets inducing $H(s,q)$-subgraphs).
Since we could not apply this improvement
in the current work, we show in 
Proposition~\ref{p:split} only the simplified version of the construction,
and give a simplified proof.
\end{remark}

\begin{corollary}
If there is an independent
$\{c;(q{-}1)n\}$-CR set in $H(n,q)$,
then there is an independent
$\{c;(q{-}1)(c{+}qn)\}$-CR set in $H(qn{+}c,q)$.
\end{corollary}

\newpage
\section[\{10,8\}-CR code in H(7,3)]{$\{10,8\}$-CR code in $H(7,3)$}
\label{s:10,8}
The first open question
for parameters of CR-$1$ codes in $H(n,3)$ was the existence of a $\{10,8\}$-CR code in $H(7,3)$.
There are no Cayley SRG over $\FF_3$
with valency~$14$. So, we searched 
for a $\{10,8\}$-CR code
in all simple Cayley graph on $\FF_3^4$
(all such graphs were enumerated by enumerating all linear
$[7,3,3]_3$ codes with
\cite{QextNewEdition}).
A CR code with quotient matrix $\left(\begin{smallmatrix}4&10\\8&6\end{smallmatrix}\right)$
was found in the coset graph $G(H)$ of the code with check matrix
$$
H = \left(\begin{array}{ccccccc}
1&0&0&0&1&1&1 \\
0&1&0&0&1&0&0 \\
0&0&1&0&0&1&0 \\
0&0&0&1&0&0&1
\end{array}\right).
$$

\section[\{21,4;2,21\}-CR code in H(11,3)]{$\{21,4;2,21\}$-CR code in $H(11,3)$}
\label{s:21,4,2,21}
There are $401$ nonisomorphic connected Cayley graphs of degree $2\cdot 11$ 
on $\FF_3^4$. Checking them all, 
we have discovered the following finding.
One (and only one) of these graphs, the coset graph $G(H)$
of the code $C(H)$ with check matrix 
$$
H = \left(\begin{array}{c@{\ \ }c@{\ \ }c@{\ \ }cc@{\ \ }c@{\ \ }c@{\ \ }cc@{\ \ }c@{\ \ }c}
1&0&0&0&0&1&1&1&1&2&2 \\
0&1&0&0&1&0&1&1&2&1&2 \\
0&0&1&0&1&1&0&1&2&2&1 \\
0&0&0&1&1&1&1&0&1&1&1
\end{array}\right),
$$
has a $\{21,4;2,21\}$-CR code
$C=\{0000, 2211, 1012, 0121, 1220, 2102\}$,
with quotient matrix 
$$\begin{pmatrix}
    1&21&0\\2&16&4\\0&21&1
\end{pmatrix}.$$
The graph has diameter $2$, spectrum 
$\{22^1,10^2,4^{12},1^{44},-5^{14},-8^8\}$;
the stabilizer of a vertex~$o$ has order $48$ and
orbits of sizes 
$1$, $8$, $8$, $6$, $12$, $8$, $24$, $12$, $2$
($8$, $8$, $6$ for neighbors of~$o$, in particular,
the code has tree coordinate orbits,
$\{0,1,2,3\}$, $\{4,5,6,7\}$, and $\{8,9,10\}$,
and the graph has tree corresponding edge orbits).
$\Aut(C)$ has order $24$, 
acts transitively on $C$ and $C^{(2)}$ and divides $C^{(1)}$ into $4$ orbits of sizes $3$, $12$, $24$, $24$.
$C^{(2)}$ is uniquely divided into two other $\{21,4;2,21\}$-CR codes $C'$ and $C''$; the partition
$(C,C',C'',C^{(1)})$ is hence equitable with quotient matrix
$$\begin{pmatrix}
    1&0&0&21\\
    0&1&0&21\\
    0&0&1&21\\
    2&2&2&16\\
\end{pmatrix}.$$

\newpage
\section[\{28;8\}-CR code in H(11,3)]{$\{28;8\}$-CR code in $H(15,3)$}\label{s:28-8}
A $\left(\begin{smallmatrix}2&28\\8&22\end{smallmatrix}\right)$-CR code
exists in the coset graphs $G$ of the codes with the following check matrices:
$$
\left(\begin{array}{ccccccccccccccc}
0&1&1&1&1&1&1&1&1&1&1&1&0&0&0\\
1&0&0&0&0&1&1&1&1&2&2&0&1&0&0\\
0&0&0&1&1&0&0&1&1&0&2&0&0&1&0\\
1&1&2&0&1&0&2&0&1&1&2&0&0&0&1\\
\end{array}\right)
$$
$$
\left(\begin{array}{ccccccccccccccc}
1&1&1&1&1&1&1&1&1&1&1&1&0&0&0\\
0&0&0&0&1&1&1&1&2&2&2&0&1&0&0\\
0&1&1&2&0&0&2&2&0&1&1&0&0&1&0\\
2&0&1&0&0&1&0&1&0&0&1&0&0&0&1\\
\end{array}\right)
$$
The inner degree of a $\{28;8\}$-CR code is $2$; 
so, it induces the union of cycles. The lengths
of cycles are $6$, $12$ or $6$, $4$, $4$, $4$
for different found codes (the search was not exhaustive);
however, the lengths of cycles for codes induced in $H(15,3)$
can be larger.

Previously, $\{28;8\}$-CR codes were known in $H(16,3)$,
from a $\{7;2\}$-CR code in $H(4,3)$ (the union of two $1$-perfect codes). It is not clear yet if there is a $\{28;8\}$-CR code (which must be an independent set) in $H(14,3)$.

\section[\{35;10\}- and \{25;20\}- CR codes in H(19,3)]{$\{35;10\}$- and $\{25;20\}$- CR codes in $H(19,3)$}\label{s:35-10}
$\left(\begin{smallmatrix}3&35\\10&28\end{smallmatrix}\right)$-CR 
and
$\left(\begin{smallmatrix}13&25\\20&18\end{smallmatrix}\right)$-CR codes
exist in the coset graph $G$ of the code with the following check matrix (as one of several?):
$$
\left(\begin{array}{c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }c@{\ \ }cc@{\ \ }cc@{\ \ }c@{\ \ }cc@{\ \ }c}
0&0&0&0&1&1&1&1&1&1&1&1 &0&0 &1&1 &0&0&0\\
1&1&1&1&0&0&0&0&1&1&1&1 &0&0 &1&0 &1&0&0\\
0&0&1&2&0&0&1&2&0&0&1&2 &1&1 &0&0 &0&1&0\\
1&2&0&0&1&2&0&0&1&2&0&0 &1&2 &0&0 &0&0&1
\end{array}\right).
$$
The graph has $81\cdot 2^4\cdot 72$ automorphisms, $4$ edge orbits
(the corresponding groups of coordinates of the code have sizes $12$, $2$, $3$, $2$).

The subgraph induced by the found $\{35;10\}$-CR code has degree $3$ (as we see from the quotient matrix) and connected components of sizes 
$6$, $4$, $4$, $4$ (so, the last three are cliques).
The code has $3456$ automorphisms and is divided into $2$ orbits of sizes $6$ and $12$ (non-code orbits have sizes $24$, $24$, $12$, $3$).

Previously, $\{35;10\}$- and $\{25;20\}$- CR codes were known in $H(20,3)$,
from $\{7;2\}$-CR and $\{5;4\}$-CR code in $H(4,3)$ (the union of two and four $1$-perfect codes, respectively). It is not clear yet if there is a $\{35;10\}$-CR code (which must be the union of independent edges) in $H(18,3)$ or a $\{25;20\}$-CR code in $H(n,3)$, $n=16,17,18$.

\newpage
\section{SRG(81,20,1,6) and SRG(81,60,45,42)}

There is only $1$ nonequivalent two-weight
$(10,4,\{6,9\})_3$ code (doubly-shortened Golay $(12,6,6)_3$ code), with check matrix
$$ H_1 = 
\begin{pmatrix}
      0&1&1&1&1&1&1&0&0&0 \\
      1&0&1&1&2&2&0&1&0&0 \\
      1&1&0&2&1&2&0&0&1&0 \\
      1&1&2&0&2&1&0&0&0&1 
\end{pmatrix}.
$$
The SRG$(81,20,1,6)$ graph $G(H_1)$ (Brouwer--Haemers graph \cite[10.28]{BvM:SRG}) has CR codes with the following i.a.:\\
eigenvalue $2$:
$\{16,2\}$ \cite{BvM:SRG}, $\{14,4\}$, $\{12,6\}$ \cite{BvM:SRG}, $\{10,8\}$;\\
eigenvalue $-7$:
$\{18;9\}$ \cite{BvM:SRG}, \underline{$\{15;12\}$} \cite{BvM:SRG}\\
covering radius~$2$:
$\{20, 18; 1, 6\}$, 
$\{18, 8; 1, 18\}$,
$\{18, 7; 2, 18\}$,
$\{18, 6; 3, 18\}$,
$\{18, 5; 4, 18\}$.

The complement, SRG$(81,60,45,42)$ has CR codes with the following i.a.:\\
eigenvalue $6$
$\{36;18\}$,
$\{30;24\}$;\\
eigenvalue $-3$:
$\{56;7\}$,
$\{49;14\}$,
$\{42;21\}$,
$\{35;28\}$;\\
covering radius~$2$:
$\{60, 14; 1, 42\}$.

We have the following new parameters of CR-$1$ codes in
$H(n,3)$:
\begin{itemize}
    \item $\{15;12\}$-CR code, $n=10$ (previous value $n=12$, lower bound $n\ge 10$).
\end{itemize}

\section{SRG(243,22,1,2)
          }

There is only $1$ SRG(243,22,1,2), the coset graph of the perfect Golay $(11,6,5)_3$ code.
It has CR codes with the following i.a.:\\
eigenvalue $4$:
$\{16,2\}$, $\{14,4\}$, $\{12,6\}$, $\{10,8\}$;\\
eigenvalue $-5$:
$\underline{\{27-c;c\}}$, $c=\underline{5,{6},{7},8},9,\underline{10,11},{12},\underline{13}$;\\
covering radius~$2$:\\
$\{18, 5; 4, 18\}$,
$\{18, 6; 3, 18\}$,
$\{18, 7; 2, 18\}$,
$\{18, 8; 1, 18\}$,
$\{20, 18; 1, 6\}$,
$\{22, 20; 1, 2\}$
(the putative arrays
$\{20, 10; 2, 13\}$,
$\{21, 10; 2, 12\}$,
$\{22, 2; 4, 17\}$,
$\{22, 9; 2, 12\}$ 
where rejected with the ILP solver CPLEX in GAMS \cite{GAMS}).


We have the following new parameters of CR-$1$ codes in
$H(n,3)$:
\begin{itemize}
    \item $\{27-c;c\}$-CR codes, $c=5,6,7,8,\,10,11,\,13$, $n=11$ (previous values $n=13,12,13,12,\,13,12,\,13$, lower bound $n\ge 11$).
\end{itemize}

\newpage
\section{DRG \{24,22,20;1,2,12\} }

The coset graph $G$ of the unique ternary
Golay $[12,6,6]_3$ code
is a distance-regular graph with
intersection array
$\{24, 22, 20; 1, 2, 12\}$ 
and eigenvalues $-12$, $-3$, $6$, $24$.

Below we list the putative intersection arrays of completely regular codes of covering radius more than~$1$ in 
$G$ meeting the following three conditions: monotonicity, eigenvalues are in $\{-12$, $-3$, $6$, $24\}$, the distance-$2$ and distance-$3$ quotient matrices are integer.

$\{18, 5; 4, 18\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{18, 6; 3, 18\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{18, 7; 2, 18\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{18, 8; 1, 18\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{20, 10; 2, 13\}$, eigenvalues: $-3$, $6$, $24$ ??? 

$\{20, 10; 8, 16\}$, eigenvalues: $-12$, $6$, $24$ ?? (GAMS: No solution)

$\{20, 18; 1, 6\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{20, 18; 4, 12\}$, eigenvalues: $-12$, $6$, $24$ ??? (GAMS: No solution)

$\{21, 10; 2, 12\}$, eigenvalues: $-3$, $6$, $24$ ???

$\{22, 2; 4, 17\}$, eigenvalues: $-3$, $6$, $24$ ???

$\{22, 9; 2, 12\}$, eigenvalues: $-3$, $6$, $24$ ???

$\{22, 20; 1, 2\}$, eigenvalues: $-3$, $6$, $24$ ++ 

$\{22, 20; 4, 8\}$, eigenvalues: $-12$, $6$, $24$ ??? (GAMS: No solution)

$\{24, 4; 2, 15\}$, eigenvalues: $-3$, $6$, $24$ ???

$\{24, 12; 6, 12\}$, eigenvalues: $-12$, $6$, $24$ ? (GAMS: $\exists$)

$\{24, 14; 1, 6\}$, eigenvalues: $-3$, $6$, $24$ ? (GAMS: $\exists$)

$\{24, 20; 4, 6\}$, eigenvalues: $-12$, $6$, $24$ ? (GAMS: No solution)

$\{24, 22, 20; 1, 2, 12\}$, eigenvalues: $-12$, $-3$, $6$, $24$  $+$

Since $G$ covers SRG$(243,22,1,2)$,
the arrays feasible in SRG$(243,22,1,2)$
also feasible in $G$ (marked as ``$++$'').

\newpage
\section{SRG(81,30,9,12) and SRG(81,50,31,30)}

There are exactly $2$ nonequivalent 
$(15,4,\{9,12\})_3$ codes \cite{HamHel96}, with check matrices as below. The corresponding SRG are discussed in~\cite[\S10.29]{BvM:SRG}.
$$
H_1 = 
\left(
\begin{array}{@{\,}c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\,}}
1&1&1&0&0&0&1&1&1&1&1&1&1&1&1\\
0&0&0&1&1&1&1&1&1&2&2&2&2&2&2\\
0&1&0&0&1&0&2&1&2&1&2&0&0&1&2\\
0&0&2&0&0&2&1&1&2&0&2&1&2&1&0 
\end{array} \right),\ 
H_2 = 
\left(
\begin{array}{@{\,}c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\,}}
1&1&1&0&0&0&1&1&1&1&1&1&1&1&1\\
0&0&0&1&1&1&1&1&1&2&2&2&2&2&2\\
0&1&0&0&1&0&2&1&2&0&1&2&0&1&2\\
0&0&2&0&0&2&1&1&2&0&2&1&2&1&0 
\end{array}  \right).
$$

$$
\begin{array}{c|c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c@{\ }c|c@{\ }c|}
\mathrm{SRG}(81,30,9,12)&\multicolumn{12}{|c|}{\mbox{eigenvalue $3$}}&
\multicolumn{2}{|c|}{\mbox{$-6$}}\\\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&
\rotatebox{90}{$\{25;2\}$} & 
\rotatebox{90}{$\{24;3\}$} & 
\rotatebox{90}{$\{23;4\}$} & 
\rotatebox{90}{$\{22;5\}$} &  
\rotatebox{90}{$\{21;6\}$} & 
\rotatebox{90}{$\{20;7\}$} & 
\rotatebox{90}{$\{19;8\}$} & 
\rotatebox{90}{$\{18;9\}$} & 
\rotatebox{90}{$\{17;10\}$} & 
\rotatebox{90}{$\{16;11\}$} & 
\rotatebox{90}{$\{15;12\}$} & 
\rotatebox{90}{$\{14;13\}$} & 
\rotatebox{90}{$\{24;12\}$} & 
\colorbox[gray]{0.9}{\rotatebox{90}{{$\{20;16\}$}}} \\\hline
116640 & + & + & + & + & + & + & + & + & + & + & + & + & + & - \\
5832 & - & + & - & - & + & - & - & + & - & - & + & - & + & + \\
\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&
\rotatebox{90}{\colorbox[gray]{0.9}{$\{50;4\}$}} & 
\rotatebox{90}{$\{48;6\}$} & 
\rotatebox{90}{\colorbox[gray]{0.9}{$\{46;8\}$}} & 
\rotatebox{90}{$\{44;10\}$} &  
\rotatebox{90}{$\{42;12\}$} & 
\rotatebox{90}{$\{40;14\}$} & 
\rotatebox{90}{$\{38;16\}$} & 
\rotatebox{90}{$\{36;18\}$} & 
\rotatebox{90}{$\{34;20\}$} & 
\rotatebox{90}{$\{32;22\}$} & 
\rotatebox{90}{$\{30;24\}$} & 
\rotatebox{90}{$\{28;26\}$} & 
\rotatebox{90}{$\{30;15\}$} & 
\rotatebox{90}{{$\{25;20\}$}} \\\hline
\mathrm{SRG}(81,50,31,30)&\multicolumn{12}{|c|}{\mbox{eigenvalue $-4$}}&
\multicolumn{2}{|c|}{\mbox{$5$}}
\end{array}
$$
CR-$2$ codes: $G(H_1)$ has $\{30,6;3,24\}$, $\{30,20;1,12\}$;
$G(H_2)$ and both complement graphs
have only trivial $\{30,20;1,12\}$-
and $\{50,18;1,30\}$-CR codes, respectively.

We have the following new parameters of CR-$1$ codes in
$H(n,3)$:
\begin{itemize}
    \item $\{20;16\}$-CR code, $n=15$ (previous value $n=16$, lower bound $n\ge 13$);
    \item $\{46;8\}$-CR code, $n=25$ (previous value $n=26$, lower bound $n\ge 23$);
    \item $\{50;4\}$-CR code, $n=25$ (previous value $n=26$, lower bound $n\ge 25$);
\end{itemize}

\newpage
\section{SRG(81,24,9,6), SRG(81,56,37,42)}
\label{s:24}

There are exactly $2$ nonequivalent ``$\FF_3$-linear'' SRG$(81,24,9,6)$, 
NVO$^+_4(3)=G(H_1)$ ($93312$ automorphisms)
and
OA$(3,\FF_9)$ ($23328$ automorphisms).
See also [W. D. Wallis, Construction of strongly regular graphs using affine designs, Bull. Austral. Math. Soc. 4 (1971) 41–49. Corrigenda, 5 (1971), p. 431. (pp. 16, 371–376,
378, 382, 385–386, 389–391, 395)]

\begin{verbatim}
H1 = [[1,1,1,1,1,1,1,1,1,0,0,0],
      [0,0,0,1,1,1,1,2,0,1,0,0],
      [0,1,1,0,0,1,1,2,0,0,1,0],
      [1,0,1,0,1,0,1,2,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H2 = [[1,1,1,1,1,1,1,1,1,0,0,0],
      [0,0,0,1,1,1,1,2,0,1,0,0],
      [0,0,1,0,1,1,2,1,0,0,1,0],
      [1,2,0,1,0,1,1,0,0,0,0,1]]
\end{verbatim}

$$
\begin{array}{c|c@{\,}c@{\,}c@{\,}c|c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c@{\,}c|}
\mathrm{SRG}(81,24,9,6)&\multicolumn{4}{|c|}{\mbox{eigenvalue $6$}}&
\multicolumn{11}{|c|}{\mbox{eigenvalue $-3$}}\\\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&
\rotatebox{90}{$\{16;2\}$} & 
\rotatebox{90}{$\{14;4\}$} & 
\rotatebox{90}{$\{12;6\}$} & 
\rotatebox{90}{$\{10;8\}$} &
\rotatebox{90}{$\{24;3\}$} &
\rotatebox{90}{\colorbox[gray]{0.9}{$\{23;4\}$}} &
\rotatebox{90}{$\{22;5\}$} &
\rotatebox{90}{$\{21;6\}$} &
\rotatebox{90}{$\{20;7\}$} &
\rotatebox{90}{$\{19;8\}$} &
\rotatebox{90}{$\{18;9\}$} &
\rotatebox{90}{$\{17;10\}$} &
\rotatebox{90}{$\{16;11\}$} &
\rotatebox{90}{$\{15;12\}$} &
\rotatebox{90}{$\{14;13\}$} 
\\\hline
93312 & - & - & + & - 
& + & + & + & + & + & + & + & +  & + & + & +   \\
23328 & + & + & + & + 
& + & - & - & + & - & - & + & -  & - & + & -   \\
\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&
\rotatebox{90}{$\{56;7\}$} & 
\rotatebox{90}{$\{49;14\}$} & 
\rotatebox{90}{$\{42;21\}$} & 
\rotatebox{90}{$\{35;28\}$} &
\rotatebox{90}{$\{48;6\}$} &
\rotatebox{90}{$\{46;8\}$} &
\rotatebox{90}{$\{44;10\}$} &
\rotatebox{90}{$\{42;12\}$} &
\rotatebox{90}{$\{40;14\}$} &
\rotatebox{90}{$\{38;16\}$} &
\rotatebox{90}{$\{36;18\}$} &
\rotatebox{90}{$\{34;20\}$} &
\rotatebox{90}{$\{32;22\}$} &
\rotatebox{90}{$\{30;24\}$} &
\rotatebox{90}{$\{28;26\}$}
 \\\hline
\mathrm{SRG}(81,56,37,42)&\multicolumn{4}{|c|}{\mbox{eigenvalue $-7$}}&
\multicolumn{11}{|c|}{\mbox{eigenvalue $2$}}
\end{array}
$$

[[18, 5, None], [None, 4, 18]]
[[18, 6, None], [None, 3, 18]]
[[18, 7, None], [None, 2, 18]]
[[18, 8, None], [None, 1, 18]]
[[20, 18, None], [None, 1, 6]]
[[21, 10, None], [None, 2, 12]]
[[22, 9, None], [None, 2, 12]]
[[24, 4, None], [None, 2, 15]]
[[24, 14, None], [None, 1, 6]]

CR-$2$ codes: $\{18,6;3,18\}$, $\{24,14;1,6\}$ in $G(H_1)$;
$\{18,6;3,18\}$, $\{24,14;1,6\}$ in 
and  $G(H_2)$;
$\{56,18;1,42\}$-CR codes in $\overline{G(H_1)}$ and  $\overline{G(H_2)}$, respectively.

We have the following new parameters of CR-$1$ codes in
$H(n,3)$:
\begin{itemize}
    \item $\{23;4\}$-CR code, $n=12$ (previous value $n=13$, lower bound $n\ge 12$).
\end{itemize}

\newpage
\section{SRG(81,32,13,12), SRG(81,48,27,30)}
\label{s:48}

There are exactly four $[16,4,\{9,12\}]_3$ codes.
The corresponding codes
have $186624$ automorphisms
(VO$^+_4(3)=G(H_1)$, see e.g.~\cite[\S3.3]{BvM:SRG}), $5184$ 
(OA$(4,\FF_9)=G(H_2)$),
$11664$, and $11664$ automorphisms, respectively.
\begin{verbatim}
H1 = [[0,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0],
      [1,0,0,0,0,1,1,1,1,1,2,2,0,1,0,0],
      [1,1,1,1,2,0,0,0,1,2,0,1,0,0,1,0],
      [1,0,1,2,0,0,1,2,1,0,1,1,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H2 = [[0,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0],
      [1,0,0,0,0,1,1,1,1,1,2,2,0,1,0,0],
      [1,0,0,1,2,0,0,1,1,2,1,1,0,0,1,0],
      [1,1,2,1,0,0,1,0,1,0,1,2,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H3 = [[0,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0],
      [1,0,0,0,0,1,1,1,1,1,2,2,0,1,0,0],
      [1,1,1,1,2,0,0,0,1,2,0,1,0,0,1,0],
      [0,0,1,2,1,0,1,2,0,1,2,2,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H4 = [[0,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0],
      [1,0,0,0,0,1,1,1,1,1,2,2,0,1,0,0],
      [1,1,1,1,2,0,0,0,1,2,0,1,0,0,1,0],
      [2,0,1,2,0,0,1,2,1,1,0,1,0,0,0,1]]
\end{verbatim}
Each of $G(H_i)$, $i=1,2,3,4$, has CR-$1$ with the following i.a.:\\
eigenvalue $5$:
$\{24;3\}$,
$\{21;6\}$,
$\{18;9\}$,
$\{15;12\}$;\\
eigenvalue $-4$:
$\{32;4\}$,
$\{28;8\}$,
$\{24;12\}$,
$\{20;16\}$;\\
covering radius~$2$: $\{32, 18; 1, 12\}$!

The complement SRG$(81, 48, 27, 30)$ have CR-$1$ with the following i.a.:\\
eigenvalue $3$:
$\{40;5\}$,
$\{35;10\}$,
$\{30;15\}$,
$\{25;20\}$;\\
eigenvalue $-6$:
$\{48;6\}$,
$\{42;12\}$,
$\{36;18\}$,
$\{30;24\}$;\\
covering radius~$2$: $\{48, 20; 1, 30\}$!

\begin{itemize}
    \item Nothing interesting found. 
    We were looking for $\{46;8\}$.
\end{itemize}

\section{SRG(81,40,19,20)}
There are exactly four $[20,4,\{12,15\}]_3$ codes; 
the corresponding graphs (with 38880, 12960, 10368, 1944 automorphisms, respectively)
are self-complementary and have CR
with the following i.a.: $\{32;4\}$, $\{28;8\}$, $\{24;12\}$, $\{20;16\}$, $\{40;5\}$, $\{35;10\}$, $\{30;15\}$, $\{25;20\}$, $\{40, 20; 1, 20\}$.

\newpage
\section{SRG(64,27,10,12), SRG(64,36,20,20)}
\label{s:27}

There is SRG$(64,27,10,12)$ VO$^-_6(2)$ ($3317760$ automorphisms).
In total, there are $5$ nonequivalent $\FF_2$-linear SRG$(64,27,10,12)$, with $3317760$, $73728$, $10240$, $7680$, and $1536$  automorphisms, respectively.

{\small
\begin{verbatim}
H1 = [[0,0,0,0,0,0,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0,],
      [0,1,1,1,1,1,0,0,0,0,0,0,0,0,0,1,1,1,1,1,1,0,1,0,0,0,0,],
      [1,0,0,0,1,1,0,0,0,0,0,1,1,1,1,0,0,1,1,1,1,0,0,1,0,0,0,],
      [1,0,0,1,0,1,0,0,1,1,1,0,0,0,1,0,1,0,1,1,1,0,0,0,1,0,0,],
      [1,0,1,0,0,1,1,1,0,0,1,0,0,1,0,0,1,1,0,1,1,0,0,0,0,1,0,],
      [1,1,0,0,0,1,0,1,0,1,1,0,1,1,1,1,0,0,0,0,1,0,0,0,0,0,1,]]
\end{verbatim}
\begin{verbatim}
H2 = [ ...,
      [1,0,0,1,0,1,0,0,1,1,1,0,0,0,1,0,1,0,1,1,1,0,0,0,1,0,0,],
      [0,0,1,0,1,1,0,1,0,1,1,0,0,1,1,1,0,0,0,1,1,0,0,0,0,1,0,],
      [0,1,0,0,1,1,1,0,1,0,1,0,1,0,1,1,0,0,1,0,1,0,0,0,0,0,1,]]
\end{verbatim}
\begin{verbatim}
H3 = [ ...,
      [0,0,0,1,1,1,0,0,1,1,1,0,0,1,1,0,1,0,0,1,1,0,0,0,1,0,0,],
      [0,0,1,1,0,1,0,1,0,0,1,0,1,1,1,1,0,0,1,0,1,0,0,0,0,1,0,],
      [1,1,0,1,0,0,1,0,0,1,0,0,1,0,1,1,0,1,1,0,1,0,0,0,0,0,1,]]
\end{verbatim}
\begin{verbatim}
H4 = [ ...,
      [1,1,0,0,1,0,1,0,0,1,1,0,0,0,1,1,0,1,0,1,1,0,0,0,0,0,1,]]
\end{verbatim}
\begin{verbatim}
H5 = [ ...,
      [1,1,0,0,1,0,1,0,0,1,0,0,1,0,1,1,1,0,0,1,1,0,0,0,0,0,1,]]
\end{verbatim}
}

$$
\begin{array}{@{\ }c|c@{\,}c@{\,}c@{\,}c|c@{\,}c@{\,}c|}
\mathrm{SRG}(64,27,10,12)&\multicolumn{4}{|c|}{\mbox{eigenvalue $3$}}&
\multicolumn{3}{|c|}{\mbox{eigenvalue~$-5$}}\\\hline
\rotatebox{0}{$|\mathrm{Aut}|$}&
\rotatebox{0}{$\{21;3\}$} & 
\rotatebox{0}{$\{18;6\}$} & 
\rotatebox{0}{$\{15;9\}$} & 
\rotatebox{0}{$\{12;12\}$} &  
\rotatebox{0}{$\{24;8\}$} & 
\rotatebox{0}{$\{20;12\}$} & 
\rotatebox{0}{$\{16;16\}$} \\\hline
3317760 & - & + & + & + & + & + & + \\
\parbox[][][l]{7em}{73728, 10240,\\ 7680, 1536} & + & + & + & + & + & + & + \\
\hline
\rotatebox{0}{$|\mathrm{Aut}|$}&
\rotatebox{0}{$\{35;5\}$} & 
\rotatebox{0}{$\{30;10\}$} & 
\rotatebox{0}{$\{25;15\}$} & 
\rotatebox{0}{$\{20;20\}$} &  
\rotatebox{0}{$\{24;8\}$} & 
\rotatebox{0}{$\{20;12\}$} & 
\rotatebox{0}{$\{16;16\}$}  
\\\hline
\mathrm{SRG}(64, 36, 20, 20)&\multicolumn{4}{|c|}{\mbox{eigenvalue $-4$}}&
\multicolumn{3}{|c|}{\mbox{eigenvalue $4$}}
\end{array}
$$

Parameters of CR-$2$ codes in each
$G(H_i)$, $i=1,2,3,4,5$:
$\{24,4;4,24\}$,
$\{24,6;2,24\}$,
$\{27,16;1,12\}$;
in each $\overline{G(H_i)}$:
$\{36, 15; 1, 20\}$.

\begin{itemize}
    \item Nothing interesting found. 
    Wanted: $\{27;5\}$, $\{21;11\}$, $\{19;13\}$, $\{17;15\}$.
\end{itemize}

\newpage
\section{SRG(64,28,12,12) and SRG(64,35,18,20)}
\label{s:28}

There are 6 $\FF_2$-linear SRG$(64,27,10,12)$, with $2580480$, $24576$, $7680$, $6144$,  $5376$, and $1536$  automorphisms, respectively.

{\footnotesize
\begin{verbatim}
H1 = [[0,0,0,0,0,0,0,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0],
      [0,0,1,1,1,1,1,0,0,0,0,0,0,0,0,0,1,1,1,1,1,1,0,1,0,0,0,0],
      [0,1,0,0,0,1,1,0,0,0,0,0,1,1,1,1,0,0,1,1,1,1,0,0,1,0,0,0],
      [1,0,0,1,1,0,0,0,0,0,1,1,0,0,1,1,1,1,0,0,1,1,0,0,0,1,0,0],
      [0,1,1,0,0,0,1,0,1,1,0,1,0,0,0,1,0,1,1,1,0,1,0,0,0,0,1,0],
      [1,0,1,0,1,1,1,1,0,1,0,0,0,1,0,1,0,0,0,1,0,1,0,0,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H2 = [[0,0,0,0,0,0,0,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,0,0,0,0,0],
      [1,1,1,1,1,1,1,0,0,0,0,0,0,0,1,1,1,1,1,1,1,1,0,1,0,0,0,0],
      [0,0,0,0,1,1,1,0,0,0,0,1,1,1,0,0,0,1,1,1,1,1,0,0,1,0,0,0],
      [0,0,0,1,0,1,1,0,0,1,1,0,0,1,0,1,1,0,0,1,1,1,0,0,0,1,0,0],
      [0,1,1,0,0,0,1,0,1,0,0,0,1,1,1,0,1,1,1,0,0,1,0,0,0,0,1,0],
      [1,0,1,0,0,0,1,1,0,0,1,0,0,1,1,1,1,0,1,0,1,0,0,0,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H3 = [...,
      [0,0,0,1,0,1,1,0,0,1,1,0,0,1,0,1,1,0,0,1,1,1,0,0,0,1,0,0],
      [0,1,1,0,0,0,1,0,1,0,0,0,1,1,1,0,1,1,1,0,0,1,0,0,0,0,1,0],
      [1,0,1,0,0,0,1,1,1,0,1,0,0,0,0,1,1,0,1,0,1,1,0,0,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H4 = [...,
      [0,0,1,1,0,0,1,0,0,1,1,0,0,1,1,1,1,0,0,0,1,1,0,0,0,1,0,0],
      [0,1,0,0,0,1,1,1,1,0,0,0,1,0,0,1,1,0,1,1,0,1,0,0,0,0,1,0],
      [1,0,0,1,0,0,1,0,1,0,1,1,1,1,0,0,1,0,0,1,0,1,0,0,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H5 = [...,
      [0,1,0,0,0,1,1,1,1,0,0,0,1,0,0,1,1,0,1,1,0,1,0,0,0,0,1,0],
      [1,1,0,1,0,0,0,0,1,0,1,0,0,1,1,0,1,1,0,1,1,0,0,0,0,0,0,1]]
\end{verbatim}
\begin{verbatim}
H6 = [...,
      [0,1,0,0,0,1,1,0,1,0,1,0,1,0,0,1,1,0,0,1,1,1,0,0,0,0,1,0],
      [1,0,0,1,1,0,0,1,0,0,0,1,1,0,1,0,1,0,1,1,0,1,0,0,0,0,0,1]]
\end{verbatim}
}

$$
\begin{array}{c|cccc|ccccccc|}
\mathrm{SRG}(64,28,12,12)&\multicolumn{4}{|c|}{\mbox{eigenvalue $4$}}&
\multicolumn{7}{|c|}{\mbox{eigenvalue $-4$}}\\\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&\rotatebox{90}{$\{21;3\}$} & 
\rotatebox{90}{$\{18;6\}$} & 
\rotatebox{90}{$\{15;9\}$} & 
\rotatebox{90}{$\{12;12\}$} &  
\rotatebox{90}{$\{28;4\}$} & 
\rotatebox{90}{$\{26;6\}$} & 
\rotatebox{90}{$\{24;8\}$} & 
\rotatebox{90}{$\{22;10\}$} & 
\rotatebox{90}{$\{20;12\}$} & 
\rotatebox{90}{$\{18;14\}$} & 
\rotatebox{90}{$\{16;16\}$} \\\hline
2580480 & + & + & + & +  & + & - & + & - & + & - & + \\
24576 & - & + & - & +  & + & + & + & + & + & + & + \\
7680 & - & + & + & +  & + & - & + & - & + & - & +  \\
6144,5376,1536  & - & + & - & +  & + & - & + & - & + & - & +  \\ 
\hline
\rotatebox{90}{$|\mathrm{Aut}|$}&\rotatebox{90}{$\{35;5\}$} & 
\rotatebox{90}{$\{30;10\}$} & 
\rotatebox{90}{$\{25;15\}$} & 
\rotatebox{90}{$\{20;20\}$} &  
\rotatebox{90}{$\{28;4\}$} & 
\rotatebox{90}{$\{26;6\}$} & 
\rotatebox{90}{$\{24;8\}$} & 
\rotatebox{90}{$\{22;10\}$} & 
\rotatebox{90}{$\{20;12\}$} & 
\rotatebox{90}{$\{18;14\}$} & 
\rotatebox{90}{$\{16;16\}$} \\\hline
\mathrm{SRG}(64,35,18,20)&\multicolumn{4}{|c|}{\mbox{eigenvalue $-5$}}&
\multicolumn{7}{|c|}{\mbox{eigenvalue $3$}}
\end{array}
$$

Parameters of CR-$2$ codes in each
$G(H_i)$, $i=1,...,6$:
$\{24,4;4,24\}$,
$\{28,15;1,12\}$;
in each $\overline{G(H_i)}$:
$\{35, 16; 1, 20\}$.

\begin{itemize}
    \item Nothing interesting found. 
    Wanted: $\{17;15\}$.
\end{itemize}

\newpage
\section[Wanted: q=3]{Wanted: $q=3$}
To find CR-$1$ codes with new parameters in $H(k/2,3)$,
we need the following Cayley graphs on $\FF_3^{m}$:
\begin{itemize}
    \item (\underline{Done, see Section~\ref{s:10,8}.}) A graph of degree $k=14$, with eigenvalue $k-3\cdot 6=-4$ and
    a $\{10;8\}$-CR code. 
    \item (\underline{Done, see Section~\ref{s:24}.}) A graph of degree $k=24$ with eigenvalue $k-3\cdot 9=-3$ and
    a $\{23;4\}$-CR code.
    \item A graph of degree $k=26$, with eigenvalue $k-3\cdot 12=-10$ and
    a $\{20;16\}$-CR code (for degree $k=28$, it is solved  by doubling the parameters in Section~\ref{s:10,8}). 
    \item A graph of degree $k=28+2s$, $s=0,1$, with eigenvalue $k-3\cdot 12=k-36$ and
    a $\{28;8\}$-CR code. \\
    (Update: found for $s=1$ in a degree-$30$ Cayley graph on $\FF_3^4$.)
    \item A graph of degree $k=32+2s$, $s=0,1,2,3$, with eigenvalue $k-3\cdot 15=k-45$ and
    a $\{25; 20\}$-CR code or
    a $\{35; 10\}$-CR code ($s=2,3$). 
    \item A graph of degree $k=40+2s$, $s=0,1$, with eigenvalue $k-3\cdot 18=k-54$ and a CR code with one of the following i.a.:
$\{28; 26\}$,
$\{32; 22\}$,
$\{34; 20\}$,
$\{38; 16\}$,
$\{40; 14\}$,
$\{30; 24\}$ ($k=-1$),
$\{42; 12\}$ ($k=1$),
$\{46; 8\}$ ($k=3$).
\end{itemize}

\section[Wanted: q=2]{Wanted: $q=2$}
To find CR-$1$ codes with new parameters in $H(k,2)$,
we need the following Cayley graphs on $\FF_2^{m}$:
\begin{itemize}
    \item A graph of degree $k=24+s$, $s=0,...,5$, with eigenvalue $k-2\cdot 16=k-32$ and a CR code with one of the following i.a.:\\
$\{27;5\}$ ($s=3$),\\
$\{25;7\}$ ($s=1,2$),\\
$\{23;9\}$ ($s=0,1,2$),\\
$\{22;10\}$ ($s=1$),\\
$\{21;11\}$ ($s=0,1,2,3$),\\
$\{19;13\}$ ($s=1,2,3$),\\
$\{17;15\}$ ($s=0,1,2,3,4,5$).\\
Update: all Cayley graphs on $\FF_2^{5}$ have been checked, nothing interesting; \url{https://sagecell.sagemath.org/?q=zecdzi}
\item A graph of degree $k=36+s$, $s=0,1,2$, with eigenvalue $k-2\cdot 24=k-48$ and a CR code with one of the following i.a.:
$\{33;15\}$.
\end{itemize}

\section
[New parameters of CR codes in H(n,q)]
{New parameters of CR codes in $H(n,q)$} \label{s:sum}

\begin{tabular}{@{\ }cccccl@{\ }}
    I.A. & $q$ & $n$($n_{\mathrm{old}}$) & l.bound & {$\tau = (b+c)/qn$} & \phantom{$= (b+c)/qn$} in graph \\[1mm] \hline
 $\{10;8\}$ & $3$ & $7(8)$ & $\ge 7$ & $6/7 \sim 0.857$ & \\[1mm]
 $\{23;4\}$ & $3$ & $12(13)$ & $\ge 12$ & $3/4 = 0.75$ &${\mathrm{VNO}^+_4(3)}$\\[1mm]
 $\{22;5\}$ & $3$ & $11(13)$ & $\ge 11$ & $9/11 \sim 0.818$ & Berlekamp--Van Lint--Seidel graph\\[1mm]
 $\{21;6\}$ & $3$ & $11(12)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}  \\[1mm]
 $\{20;7\}$ & $3$ & $11(13)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}   \\[1mm]
 $\{19;8\}$ & $3$ & $11(12)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}   \\[1mm]
 $\{17;10\}$ & $3$ & $11(13)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}   \\[1mm]
 $\{16;11\}$ & $3$ & $11(12)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}   \\[1mm]
 $\{15;12\}$ & $3$ & $10(12)$ & $\ge 10$ & $9/10=0.9$ & Brouwer--Haemers graph \cite[10.28]{BvM:SRG}\\[1mm]
 $\{14;13\}$ & $3$ & $11(13)$ & $\ge 11$ & $9/11 \sim 0.818$ & B.--v.L.--S. \cite[10.55]{BvM:SRG}   \\[1mm]
 $\{20;16\}$ & $3$ & $15(16)$ & $\ge 13$ & $4/5 =0.8$ & $\Delta\ne{\mathrm{VNO}^-_4(3)}$ \cite[10.29]{BvM:SRG}\\[1mm]
   &   & $14(16)$ & $\ge 13$ & $6/7 \sim 0.857$ & $2\times\{10;8\}$ \\[1mm]
 $\{28;8\}$ & $3$ & $15(16)$ & $\ge 14$ & $4/5 =0.8$ & $\Cay(\FF_3^4)$, Section~\ref{s:28-8}\\[1mm]
 $\{35;10\}$ & $3$ & $19(20)$ & $\ge 18$ & $15/19 \sim 0.789$ & $\Cay(\FF_3^4)$, Section~\ref{s:35-10}\\[1mm]
 $\{25;20\}$ & $3$ & $19(20)$ & $\ge 16$ & $15/19 \sim 0.789$ & $\Cay(\FF_3^4)$, Section~\ref{s:35-10}\\[1mm]
 $\{46;8\}$ & $3$ & $25(26)$ & $\ge 23$ & $18/25=0.72$ & $\overline{\mathrm{VNO}^-_4(3)}$ \cite[10.29]{BvM:SRG}\\[1mm]
   &  & $24(26)$ & $\ge 23$ & $3/4=0.75$ & $2\times\{23;4\}$  \\[1mm]
 $\{50;4\}$ & $3$ & $25(26)$ & $\ge 25$ & $18/25=0.72$ & $\overline{\mathrm{VNO}^-_4(3)}$  \cite[10.29]{BvM:SRG}\\[1mm]
 $\{21,4;2,21\}$ & $3$ & $11(-)$ & $\ge 11$ & --- & $\Cay(\FF_3^4)$, Section~\ref{s:21,4,2,21} 
\end{tabular}

We have searched:
Cayley graphs on $\FF_3^4$ up to degree~$54$,
Cayley graphs on $\FF_2^6$ up to degree~$21$,
and Cayley graphs on $\FF_2^6$ of degree~$22$ 
with at least $2^6\cdot 16$ automorphisms.

\newpage

\nocite{HamHel96}
\nocite{BFWW2006}
\nocite{BvM:SRG}
\nocite{CRCinDRG}
\nocite{KooKroMar:Ch7tables}

\providecommand\href[2]{#2} \providecommand\url[1]{\href{#1}{#1}}
  \def\DOI#1{{\href{https://doi.org/#1}{https://doi.org/#1}}}\def\DOIURL#1#2{{\href{https://doi.org/#2}{https://doi.org/#1}}}

\end{document}